# Higher-weight Jacobians

### v4

@@ 3.3. Computing cohomology of complex tori
The following discussion is based on Chapter I of [mumabvar].
@@
The natural inclusion of the constant sheaf $\underline{\mathbf{Z}}$ in the structure sheaf $\mathscr{O}_{X}$ gives rise to a map of the cohomology groups ${\rm H}^{r}(X,\mathbf{Z})\to{\rm H}^{r}(X,\mathscr{O}_{X})$; it can be shown that this arises as the exterior power of the map
@@
In the setting of Lemma 3.3.4, the lattice $\mathfrak{L}_{m}(X)$ for $m>2$ is homothetic to the lattice If $X=E_{1}\times\dots\times E_{n}$ for isogenous CM elliptic curves $E_{i}=\mathbf{C}/\Lambda_{i}$, the lattice $\mathfrak{L}(X)$ is homothetic to the lattice
@@ 4.1. Computing the Brauer-Jacobian with class groups
There is a bijective correspondence between the set of isomorphism classes of elliptic curves with CM by $R$ and $\operatorname{Cl}(R)$. Namely, if our elliptic curve $E$ is $\mathbf{C}/\Lambda$ for the lattice $\Lambda=\langle 1,\tau\rangle$ and ${\rm End}(E)={\rm End}(\Lambda)=R$, then $\Lambda\subseteq{\rm Frac}(R)$ must be an invertible (proper) fractional ideal of $R$. We will write $[E]$ or $[\Lambda]$ for the element of $\operatorname{Cl}(R)$ associated to $E=\mathbf{C}/\Lambda$ via this correspondence.
@@
For two CM elliptic curves $E_{1},E_{2}$ with ${\rm End}(E_{1})={\rm End}(E_{2})=R$, the ring $R$ is an order in an imaginary quadratic number field $K$ with class group $\operatorname{Cl}(R)$, and $E_{i}=\mathbf{C}/\Lambda_{i}$ with $\Lambda_{i}$ a proper fractional ideal of $R$. Then $\mathfrak{C}(E_{1}\times E_{2})$ is a complex elliptic curve such that
$[E_{1}][E_{2}]=[\mathfrak{C}(E_{1}\times E_{2})]\in\operatorname{Cl}(R).$
@@
By Lemma 3.3.2 we know that $\mathfrak{L}(\Lambda_{1}\times\Lambda_{2})$ is homothetic to $\Lambda_{1}\Lambda_{2}$. By Lemma 3.1.4 it is clear that $[\Lambda_{1}][\Lambda_{2}]=[\Lambda_{1}\Lambda_{2}]=[\mathfrak{L}(\Lambda_{1}\times\Lambda_{2})]\in\operatorname{Cl}(R)$, from which it follows that $[\mathfrak{C}(E_{1}\times E_{2})]=[E_{1}][E_{2}]\in\operatorname{Cl}(R)$. ∎
@@
$[\mathfrak{C}(E\times E)]=[E]^{2}\in\operatorname{Cl}({\rm End}(E)).$
@@
Thus, in the particular case where ${\rm End}(E_{1})={\rm End}(E_{2})$ (so $a=b=c$), we have $[E_{1}][E_{2}]=[\mathfrak{C}(E_{1}\times E_{2})]\in\operatorname{Cl}(\mathbf{Z}[a\alpha]).$
@@
$[\Lambda_{1}\dots\Lambda_{n}]=\prod_{i=1}^{n}\phi_{d,f_{i},\alpha}([E_{i}])\in\operatorname{Cl}(\mathcal{O}_{0})$
@@ 4.2. Which complex tori have Brauer-Jacobians?
We have been able to describe the Brauer-Jacobian of a product of pairwise isogenous CM elliptic curves in the previous sections. It has been shown ([sm74] for the case of surfaces, [schoen] for the general case) that a complex abelian variety which is isogenous to a product of isogenous CM elliptic curves must be isomorphic to such a product. In the following theorem we use this fact to show that among complex tori, only such products of elliptic curves have Brauer-Jacobians.
@@
It is immediately clear that (1) implies (2); the reverse implication is given by [schoen, Satz 2.4]. It is also immediate that (4) implies (3).
@@
By [mumabvar, Chapter I.2, Theorem of Appell-Humbert], this kernel is isomorphic to the Neron-Severi group ${\rm NS}(X)\simeq\rm{Pic}(X)/\rm{Pic}^{0}(X)$, which is always a free abelian group of finite rank (see either [blabvar, Corollary 5.3.9] or [mumabvar, Chapter IV.19, Corollary 2, Theorem 3]). Furthermore, the rank of the Neron-Severi group is $n^{2}$ exactly in the situation of (2) [blabvar, Proposition 5.2.1, Corollary 5.3.8]. Therefore (3) implies (2).
Then (2) implies (1) implies (4) implies (3) implies (2). By [blabvar, Proposition 5.2.1, Corollary 5.3.8], (5) is equivalent to (2). Thus, all conditions are equivalent, as desired. ∎
@@
We note that for any complex variety $X$, the cokernel of the map ${\rm H}^{2}(X,\mathbf{Z})\to{\rm H}^{2}(X,\mathscr{O})$ is a complex torus when the Neron-Severi group ${\rm NS}(X)$ is of ‘maximal’ rank; in [beauville] this is referred to as $X$ being $\rho$-maximal. In particular a complex abelian variety is $\rho$-maximal if and only if it satisfies the equivalent conditions of Theorem 1.2.
@@
Similarly to the term $\rho$-maximal of [beauville], we say a complex variety $X$ is $q$-maximal if the cokernel of the map ${\rm H}^{2}(X,\mathbf{Z})\to{\rm H}^{2}(X,\mathscr{O})$ is a complex torus, so $X$ has a $q$-Jacobian. We see that $X$ is $2$-maximal if and only if it is $\rho$-maximal if and only if it has a Brauer-Jacobian.
The proof in [schoen, Satz 2.4] of the fact that any complex abelian variety which is isogenous to $E^{n}$ for a CM elliptic curve $E$ is in fact isomorphic to a product of elliptic curves is ‘algebraic’, as it is a consequence of [bass, Theorem 1.7], which states that for a certain class of rings (to which all orders in imaginary quadratic number fields belong) any torsion-free module can be decomposed into a direct sum of rank $1$ modules. However, this decomposition is not unique; it follows that for $X\simeq E_{1}\times\dots\times E_{n}$, the $E_{i}$ are also not unique. This ambiguity of decomposition is discussed in [sm74] for the case $n=2$ using binary quadratic forms, which we restate here. .
@@
    and for $E_{i}=\mathbf{C}/\Lambda_{i},i=1,\dots,4$ pairwise isogenous CM elliptic curves with CM by the imaginary quadratic field $K$, we have $E_{1}\times E_{2}\simeq E_{3}\times E_{4}$ if and only if ${\rm End}(E_{1})\cap{\rm End}(E_{2})={\rm End}(E_{3})\cap{\rm End}(E_{4})$ and $\Lambda_{1}\Lambda_{2}=\Lambda_{3}\Lambda_{4}$.
@@
Due to [schoen] (and similarly to [bf68, Lemma 8]) we see that the lattices $\Lambda_{1}\times\Lambda_{2}\subseteq\mathbf{C}^{2}$ and $\mathcal{O}(\mathcal{A})\times\Lambda_{1}\Lambda_{2}$ are isomorphic to one another as $\mathcal{O}(\mathcal{A})$-modules, which proves the main statement of the corollary, from which (i) follows.
@@
We can also use Theorem 4.2.3 to apply the results of Section 3.2 concerning the ‘field of definition’ of elliptic curves to the question of when a product $\mathcal{A}=E_{1}\times E_{2}$ of elliptic curves with CM by an order $\mathcal{O}$ and its Brauer-Jacobian $\mathfrak{C}(\mathcal{A})$ can both be defined over a field ‘smaller’ than the ring class field $L_{\mathcal{O}}$ (in the primitive case).
@@
It follows from Theorem 4.2.3(ii) that the condition that ${\rm End}(\mathfrak{C}(\mathcal{A}))=\mathcal{O}(\mathcal{A})$ for $\mathcal{A}$ a $2$-maximal complex abelian surface is equivalent to the condition that for any $E_{1},E_{2}$ with $\mathcal{A}\simeq E_{1}\times E_{2}$ we have ${\rm End}(E_{1})={\rm End}(E_{2})=\mathcal{O}(\mathcal{A})$. We set $\mathcal{O}=\mathcal{O}(\mathcal{A})$ for ease of writing in this proof. In fact, for any $E_{1},E_{2}$ with CM by $\mathcal{O}$, we have $\mathcal{A}\simeq E_{1}\times E_{2}$ if and only if $[E_{1}][E_{2}]=[\mathfrak{C}(\mathcal{A})]\in\operatorname{Cl}(\mathcal{O})$ by Theorem 4.2.3 and Theorem 4.1.4. We can then apply Lemma 3.2.2 to see that $\mathbf{Q}(j(E_{1}))=\mathbf{Q}(j(E_{2}))=\mathbf{Q}(j(\mathfrak{C}(\mathcal{A})))$ if and only if $[E_{1}][E_{2}]^{-1},[E_{1}],$ and $[E_{2}]$ all have order $\leq 2$. Since $\operatorname{Cl}(\mathcal{O})$ is abelian, this implies that $[\mathfrak{C}(\mathcal{A})]=[E_{1}][E_{2}]$ has order $\leq 2$ also. This completes one direction of the proof.
If $[\mathfrak{C}(\mathcal{A})]$ has order $\leq 2$, then $\mathbf{Q}(j(\mathfrak{C}(\mathcal{A})))=\mathbf{Q}(j(\mathbf{C}/\mathcal{O}))$, so we can set $E_{1}=\mathfrak{C}(\mathcal{A}),E_{2}=\mathbf{C}/\mathcal{O}$ to get the desired decomposition.∎
@@
Note that in Proposition 1.2, the group element $[\mathfrak{C}(\mathcal{A})]\in\operatorname{Cl}(\mathcal{O})$ has order $\leq 2$ if and only if $j(\mathfrak{C}(\mathcal{A}))$ is real. (This is true since one can show from the definition of the $j$-invariant of a lattice that $j(L)=\overline{j(\overline{L})}$, and if $L$ is a proper fractional ideal of an order $\mathcal{O}$, we have $[L][\overline{L}]=[\mathcal{O}]=1\in\operatorname{Cl}(\mathcal{O})$. See [cox11, Theorem 10.9, Exercise 11.1].)
@@
Recall from Example 4.1.3 that the order $\mathfrak{R}^{\prime}=\mathbf{Z}[\sqrt{-36}]$ in the field $K=\mathbf{Q}(i)$ has ring class field $L_{\mathfrak{R}}^{\prime}=K(\sqrt[4]{12})$ and class group $\operatorname{Cl}(\mathfrak{R}^{\prime})=\mathbf{Z}/4\mathbf{Z}$, with the identity being $1=[\mathfrak{R}^{\prime}]$, and generator $[\langle 3,1+2i\rangle]$; we can compute $[\langle 3,1+2i\rangle]^{2}=[\langle 3,2i\rangle]$.
@@
However, by Theorem 4.2.3(ii), we see that $\mathcal{A}\cong\mathbf{C}/\mathfrak{R}^{\prime}\times\mathbf{C}/\langle 3,2i\rangle$, and $\mathbf{Q}(j(\mathfrak{R}^{\prime}))=\mathbf{Q}(j(\langle 3,2i\rangle))=\mathbf{Q}[\sqrt[4]{12}]$. Therefore $\mathcal{A}$ can be defined over the minimal field of definition of $\mathfrak{C}(\mathcal{A})$, as we would expect from Proposition 1.2 since the order of $[\mathfrak{C}(\mathcal{A})]=[\langle 3,2i\rangle]$ is $2$. ∎
@@
Recall from Example 3.2.3 that the order $\mathfrak{O}=\mathbf{Z}[\sqrt{-27}]$ in the field $\mathbf{Q}(\zeta_{3})$ has ring class field $L_{\mathfrak{O}}=\mathbf{Q}(\zeta_{3},\sqrt[3]{2})$ and class group $\operatorname{Cl}(\mathfrak{O})=\mathbf{Z}/3\mathbf{Z}$, with the identity being $1=[\mathfrak{O}]$, and generator $[\Lambda_{1}]=[\langle 3,2+\sqrt{-3}\rangle]$; we can compute $[\langle 3,2+\sqrt{-3}\rangle]^{2}=[\langle 3,1+\sqrt{-3}\rangle]=[\Lambda_{2}]$.
@@
By Theorem 4.2.3(ii), we see that the only other way to decompose $\mathcal{A}^{\prime}$ as a product of two elliptic curves is as $\mathcal{A}^{\prime}=(\mathbf{C}/\Lambda_{1})^{2}\cong\mathbf{C}/\mathfrak{O}\times\mathbf{C}/\Lambda_{2},$ but $j(\mathfrak{O})\notin=\mathbf{Q}(j(\Lambda_{2}))=\mathbf{Q}(\zeta_{3}^{2}\sqrt[3]{2})$—as we would expect from Proposition 1.2, since the order of $[\mathfrak{C}(\mathcal{A}^{\prime})]=[\Lambda_{2}]$ is $3$, we can’t define $\mathcal{A}^{\prime}$ over the minimal field of definition of $\mathfrak{C}(\mathcal{A}^{\prime})$. ∎

### v5

@@ 3.3. Computing cohomology of complex tori
The following discussion is based on Chapter I of [MRM74].
@@
The natural inclusion of the constant sheaf $\underline{\mathbf{Z}}$ in the structure sheaf $\mathscr{O}_{X}$ gives rise to a map of the cohomology groups ${\rm H}^{r}(X,\mathbf{Z})\to{\rm H}^{r}(X,\mathscr{O}_{X})$; this map is the exterior power $\bigwedge^{r}$ of the map
@@
In the setting of Lemma 3.3.4, the lattice $\mathfrak{L}_{m}(X)$ for $m>2$ is homothetic to the lattice
@@ 4.1. Computing the Brauer-Jacobian with class groups
There is a bijective correspondence between the set of isomorphism classes of elliptic curves with CM by $R$ and $\operatorname{Cl}(R)$. Namely, if our elliptic curve $E$ is $\mathbf{C}/\Lambda$ for the lattice $\Lambda=\langle 1,\tau\rangle$ and ${\rm End}(E)={\rm End}(\Lambda)=R$, then $\Lambda\subseteq{\rm Frac}(R)$ must be an invertible (proper) fractional ideal of $R$. We will write $[E]$ or $\Lambda]$ for the element of $\operatorname{Cl}(R)$ associated to $E=\mathbf{C}/\Lambda$ via this correspondence.
@@
Let $R$ be an order in an imaginary quadratic number field $K$ with class group $\operatorname{Cl}(R)$. For $\Lambda_{1},\Lambda_{2}$ proper fractional ideals of $R$, the quotients $E_{i}=\mathbf{C}/\Lambda_{i}$ are CM elliptic curves with ${\rm End}(E_{1})={\rm End}(E_{2})=R$. Then $\mathfrak{C}(E_{1}\times E_{2})$ is a complex elliptic curve such that
$E_{1}][E_{2}]=[\mathfrak{C}(E_{1}\times E_{2})]\in\operatorname{Cl}(R).$
@@
By Lemma 3.3.2 we know that $\mathfrak{L}(\Lambda_{1}\times\Lambda_{2})$ is homothetic to $\Lambda_{1}\Lambda_{2}$. By Lemma 3.1.4 it is clear that $\Lambda_{1}][\Lambda_{2}]=[\Lambda_{1}\Lambda_{2}]=[\mathfrak{L}(\Lambda_{1}\times\Lambda_{2})]\in\operatorname{Cl}(R)$, from which it follows that $\mathfrak{C}(E_{1}\times E_{2})]=[E_{1}][E_{2}]\in\operatorname{Cl}(R)$. ∎
@@
$\mathfrak{C}(E\times E)]=[E]^{2}\in\operatorname{Cl}({\rm End}(E)).$
@@
Thus, in the particular case where ${\rm End}(E_{1})={\rm End}(E_{2})$ (so $a=b=c$), we have $E_{1}][E_{2}]=[\mathfrak{C}(E_{1}\times E_{2})]\in\operatorname{Cl}(\mathbf{Z}[a\alpha]).$
@@
$\Lambda_{1}\dots\Lambda_{n}]=\prod_{i=1}^{n}\phi_{d,f_{i},\alpha}([E_{i}])\in\operatorname{Cl}(\mathcal{O}_{0})$
@@ 4.2. Which complex tori have Brauer-Jacobians?
We have been able to describe the Brauer-Jacobian of a product of pairwise isogenous CM elliptic curves in the previous sections. It has been shown ([SM74] for the case of surfaces, [Sch92] for the general case) that a complex abelian variety which is isogenous to a product of isogenous CM elliptic curves must be isomorphic to such a product. In the following theorem we use this fact to show that among complex tori, only such products of elliptic curves have Brauer-Jacobians.
@@
It is immediately clear that (1) implies (2); the reverse implication is given by [Sch92, Satz 2.4]. It is also immediate that (4) implies (3).
@@
By [MRM74, Chapter I.2, Theorem of Appell-Humbert], this kernel is isomorphic to the Neron-Severi group ${\rm NS}(X)\simeq\rm{Pic}(X)/\rm{Pic}^{0}(X)$, which is always a free abelian group of finite rank (see either [BL04, Corollary 5.3.9] or [MRM74, Chapter IV.19, Corollary 2, Theorem 3]). Furthermore, the rank of the Neron-Severi group is $n^{2}$ exactly in the situation of (2) [BL04, Proposition 5.2.1, Corollary 5.3.8]. Therefore (3) implies (2).
Then (2) implies (1) implies (4) implies (3) implies (2). By [BL04, Proposition 5.2.1, Corollary 5.3.8], (5) is equivalent to (2). Thus, all conditions are equivalent, as desired. ∎
@@
We note that for any complex variety $X$, the cokernel of the map ${\rm H}^{2}(X,\mathbf{Z})\to{\rm H}^{2}(X,\mathscr{O})$ is a complex torus when the Neron-Severi group ${\rm NS}(X)$ is of ‘maximal’ rank; in [Bea14] this is referred to as $X$ being $\rho$-maximal. In particular a complex abelian variety is $\rho$-maximal if and only if it satisfies the equivalent conditions of Theorem 1.2.
@@
Similarly to the term $\rho$-maximal of [Bea14], we say a complex variety $X$ is $q$-maximal if the cokernel of the map ${\rm H}^{2}(X,\mathbf{Z})\to{\rm H}^{2}(X,\mathscr{O})$ is a complex torus, so $X$ has a $q$-Jacobian. We see that $X$ is $2$-maximal if and only if it is $\rho$-maximal if and only if it has a Brauer-Jacobian.
The proof in [Sch92, Satz 2.4] of the fact that any complex abelian variety which is isogenous to $E^{n}$ for a CM elliptic curve $E$ is in fact isomorphic to a product of elliptic curves is ‘algebraic’, as it is a consequence of [Bas62, Theorem 1.7], which states that for a certain class of rings (to which all orders in imaginary quadratic number fields belong) any torsion-free module can be decomposed into a direct sum of rank $1$ modules. However, this decomposition is not unique; it follows that for $X\simeq E_{1}\times\dots\times E_{n}$, the $E_{i}$ are also not unique. This ambiguity of decomposition is discussed in [SM74] for the case $n=2$ using binary quadratic forms, which we restate here. .
@@
    and for $E_{i}=\mathbf{C}/\Lambda_{i},i=1,\dots,4$ pairwise isogenous CM elliptic curves with CM by the imaginary quadratic field $K$, we have $E_{1}\times E_{2}\simeq E_{3}\times E_{4}$ if and only if ${\rm End}(E_{1})\cap{\rm End}(E_{2})={\rm End}(E_{3})\cap{\rm End}(E_{4})$ and $\Lambda_{1}\Lambda_{2}$ is homothetic to $\Lambda_{3}\Lambda_{4}$.
@@
Due to [Sch92] (and similarly to [BF68, Lemma 8]) we see that the lattices $\Lambda_{1}\times\Lambda_{2}\subseteq\mathbf{C}^{2}$ and $\mathcal{O}(\mathcal{A})\times\Lambda_{1}\Lambda_{2}$ are isomorphic to one another as $\mathcal{O}(\mathcal{A})$-modules, which proves the main statement of the corollary, from which (i) follows.
@@
We can also use Theorem 4.2.3 to apply the results of Section 3.2 concerning the field of definition of elliptic curves to the question of when a product $\mathcal{A}=E_{1}\times E_{2}$ of elliptic curves with CM by an order $\mathcal{O}$ and its Brauer-Jacobian $\mathfrak{C}(\mathcal{A})$ can both be defined over a field smaller than the ring class field $L_{\mathcal{O}}$ (in the primitive case).
@@
It follows from Theorem 4.2.3(ii) that the condition that ${\rm End}(\mathfrak{C}(\mathcal{A}))=\mathcal{O}(\mathcal{A})$ for $\mathcal{A}$ a $2$-maximal complex abelian surface is equivalent to the condition that for any $E_{1},E_{2}$ with $\mathcal{A}\simeq E_{1}\times E_{2}$ we have ${\rm End}(E_{1})={\rm End}(E_{2})=\mathcal{O}(\mathcal{A})$. We set $\mathcal{O}=\mathcal{O}(\mathcal{A})$ for ease of writing in this proof. In fact, for any $E_{1},E_{2}$ with CM by $\mathcal{O}$, we have $\mathcal{A}\simeq E_{1}\times E_{2}$ if and only if $E_{1}][E_{2}]=[\mathfrak{C}(\mathcal{A})]\in\operatorname{Cl}(\mathcal{O})$ by Theorem 4.2.3 and Theorem 4.1.4. We can then apply Lemma 3.2.2 to see that $\mathbf{Q}(j(E_{1}))=\mathbf{Q}(j(E_{2}))=\mathbf{Q}(j(\mathfrak{C}(\mathcal{A})))$ if and only if $E_{1}][E_{2}]^{-1},[E_{1}],$ and $E_{2}]$ all have order $\leq 2$. Since $\operatorname{Cl}(\mathcal{O})$ is abelian, this implies that $\mathfrak{C}(\mathcal{A})]=[E_{1}][E_{2}]$ has order $\leq 2$ also. This completes one direction of the proof.
If $\mathfrak{C}(\mathcal{A})]$ has order $\leq 2$, then $\mathbf{Q}(j(\mathfrak{C}(\mathcal{A})))=\mathbf{Q}(j(\mathbf{C}/\mathcal{O}))$, so we can set $E_{1}=\mathfrak{C}(\mathcal{A}),E_{2}=\mathbf{C}/\mathcal{O}$ to get the desired decomposition.∎
@@
Note that in Proposition 1.2, the group element $\mathfrak{C}(\mathcal{A})]\in\operatorname{Cl}(\mathcal{O})$ has order $\leq 2$ if and only if $j(\mathfrak{C}(\mathcal{A}))$ is real. (This is true since one can show from the definition of the $j$-invariant of a lattice that $j(L)=\overline{j(\overline{L})}$, and if $L$ is a proper fractional ideal of an order $\mathcal{O}$, we have $L][\overline{L}]=[\mathcal{O}]=1\in\operatorname{Cl}(\mathcal{O})$. See [Cox11, Theorem 10.9, Exercise 11.1].)
@@
Recall from Example 4.1.3 that the order $\mathfrak{R}^{\prime}=\mathbf{Z}[\sqrt{-36}]$ in the field $K=\mathbf{Q}(i)$ has ring class field $L_{\mathfrak{R}}^{\prime}=K(\sqrt[4]{12})$ and class group $\operatorname{Cl}(\mathfrak{R}^{\prime})=\mathbf{Z}/4\mathbf{Z}$, with the identity being $1=[\mathfrak{R}^{\prime}]$, and generator $\langle 3,1+2i\rangle]$; we can compute $\langle 3,1+2i\rangle]^{2}=[\langle 3,2i\rangle]$.
@@
However, by Theorem 4.2.3(ii), we see that $\mathcal{A}\cong\mathbf{C}/\mathfrak{R}^{\prime}\times\mathbf{C}/\langle 3,2i\rangle$, and $\mathbf{Q}(j(\mathfrak{R}^{\prime}))=\mathbf{Q}(j(\langle 3,2i\rangle))=\mathbf{Q}[\sqrt[4]{12}]$. Therefore $\mathcal{A}$ can be defined over the minimal field of definition of $\mathfrak{C}(\mathcal{A})$, as we would expect from Proposition 1.2 since the order of $\mathfrak{C}(\mathcal{A})]=[\langle 3,2i\rangle]$ is $2$. ∎
@@
Recall from Example 3.2.3 that the order $\mathfrak{O}=\mathbf{Z}[\sqrt{-27}]$ in the field $\mathbf{Q}(\zeta_{3})$ has ring class field $L_{\mathfrak{O}}=\mathbf{Q}(\zeta_{3},\sqrt[3]{2})$ and class group $\operatorname{Cl}(\mathfrak{O})=\mathbf{Z}/3\mathbf{Z}$, with the identity being $1=[\mathfrak{O}]$, and generator $\Lambda_{1}]=[\langle 3,2+\sqrt{-3}\rangle]$; we can compute $\langle 3,2+\sqrt{-3}\rangle]^{2}=[\langle 3,1+\sqrt{-3}\rangle]=[\Lambda_{2}]$.
@@
By Theorem 4.2.3(ii), we see that the only other way to decompose $\mathcal{A}^{\prime}$ as a product of two elliptic curves is as $\mathcal{A}^{\prime}=(\mathbf{C}/\Lambda_{1})^{2}\cong\mathbf{C}/\mathfrak{O}\times\mathbf{C}/\Lambda_{2},$ but $j(\mathfrak{O})\notin=\mathbf{Q}(j(\Lambda_{2}))=\mathbf{Q}(\zeta_{3}^{2}\sqrt[3]{2})$—as we would expect from Proposition 1.2, since the order of $\mathfrak{C}(\mathcal{A}^{\prime})]=[\Lambda_{2}]$ is $3$, we can’t define $\mathcal{A}^{\prime}$ over the minimal field of definition of $\mathfrak{C}(\mathcal{A}^{\prime})$. ∎